OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (shamim):

required equation is (x-h)^2+(y-k)^2=a^2 given h=3 k=2 so ur equation will b (x-3)^2+(y-2)^2=a^2

OpenStudy (shamim):

now we hv to find out the value of rdius a

OpenStudy (shamim):

radius a will b the distance between centre n ur given point